# Multiplexing in Networks and Diffusion

### v1

@@ Abstract.
Social and economic networks are often multiplexed, meaning that people are connected by different types of relationships—such as borrowing goods and giving advice. We make three contributions to the study of multiplexing. First, we document empirical multiplexing patterns in Indian village data: relationships such as socializing, advising, helping, and lending are correlated but distinct, while commonly used proxies for networks based on ethnicity and geography are nearly uncorrelated with actual relationships. Second, we examine how these layers and their overlap affect information diffusion in a field experiment. The advice network is the best predictor of diffusion, but combining layers improves predictions further. Villages with greater overlap between layers (more multiplexing) experience less overall diffusion. This leads to our third contribution: developing a model and theoretical results about diffusion in multiplex networks. Multiplexing slows the spread of simple contagions, such as diseases or basic information, but can either impede or enhance the spread of complex contagions, such as new technologies, depending on their virality. Finally, we identify differences in multiplexing by gender and connectedness. These have implications for inequality in diffusion-mediated outcomes such as access to information and adherence to norms.
@@ 1. Introduction
People maintain many different types of relationships—for example, collaborating with colleagues at work, relying on friends and acquaintances for assistance and advice, and engaging in the borrowing and lending of money or goods with family and friends. A given pair of people can have multiple such relationships. For example, college students’ partners in social activities overlap with, but also differ from, the people to whom they turn during times of stress or for academic collaboration (Morelli et al., 2017; Jackson et al., 2024).
@@
The coexistence of distinct types of relationships among the same population is known as multiplexing (see, e.g., Kivela et al. (2014)), and the interdependence of different types of relationships has been discussed since Simmel (1908). Although numerous case studies have examined multiple relationships, many basic questions remain open concerning the patterns of multiplexing in social and economic relationships, as well as how multiplexing impacts outcomes of interest such as information diffusion.
In this paper, we make three contributions.
First, in Section 2, we perform unsupervised statistical analyses on correlation structures across network layers in two large datasets. We document both significant correlation between different network layers (types of relationships) and meaningful differences in their patterns. We show that layers of informational relationships, financial relationships, and social relationships, among others, exhibit strong correlations in a sample of 143 villages in Karnataka, India, comprising nearly 30,000 households (Banerjee et al., 2013, 2019; Banerjee et al., 2024b). At the same time, different layers display distinct patterns and differ in density and other network statistics. We also show that proxies for social relationships that are commonly used in the peer effects literature, such as geographic proximity or co-ethnicity (in our data, being members of the same jati, or subcaste), are nearly orthogonal to the other layers. This suggests that relational variables constructed based on geographic or ethnic covariates can fail to serve as accurate surrogates for actual social and economic relationships.
Second, we use supervised statistical methods to show that distinctions among layers are substantively important for the study of economic outcomes, specifically the diffusion of information or behaviors. While one might have expected little difference in the predictive power of different network layers for outcomes of interest, we find that the different layers contain distinct information and combine to form a nuanced overall picture. Using data from a randomized controlled trial of information diffusion, we show that some layers are more predictive of diffusion than others—with an “advice” layer standing out—and moreover that using a suitable combination of layers yields predictions significantly better than those based on any single elicited layer. A combination of layers also affords better predictions than using the union or intersection of layers. These findings indicate that the elicited layers are not simply noisy observations of a latent one-dimensional relationship, instead containing information richer than any one-dimensional summary. Without properly accounting for the multiplexed nature of relationships, researchers may arrive at misleading conclusions about peer effects and influence.
An additional finding regarding ethnic links further supports the point that links are most usefully viewed as multi-dimensional. We show that, while the jati layer is the least predictive of diffusion and not a good proxy for actual relationships, combining it with other layers significantly improves diffusion predictions. Thus, although jati is not a good substitute for elicited network data, [^1] As we discuss below, using the jati variable drastically over-predicts links within jati, and under-predicts them across jati. One conjecture as to why the jati layer helps in predicting diffusion is that patterns of information passing on the network are related to jati. it can serve as a valuable complement.
We close Section 2 with an important and novel empirical observation: villages that are more multiplexed (have more strongly correlated layers) experience less diffusion. This correlation between layer overlap and diffusion sets the stage for our theoretical analysis.
@@
Motivated by the fact just presented, in Section 3 we develop our third main contribution—a new model of and theoretical results on how multiplexing relates to diffusion. First, we model how the degree of multiplexing affects a standard diffusion or contagion process in which a person may be infected/informed by any single infected other (called simple contagion). We introduce a definition capturing what it means for an individual to be unambiguously more multiplexed in one multiplex network than another. When such a comparison can be made, we prove that the more multiplexed individual is less likely to become infected for any given probability of neighbor infection. Building on this result, we demonstrate that in a standard SIS (Susceptible-Infected-Susceptible) simple contagion model, the steady-state infection rate decreases as individuals become more multiplexed. These results can be summarized by saying that multiplexing impedes simple diffusions. We then develop a theory of how multiplexing impacts complex diffusion processes—ones in which people only become infected or adopt a new behavior/practice if they experience sufficiently many interactions with infected others. Here we show that multiplexing can either enhance or impede diffusion, depending on the virality of the process. The nonmonotonicities identified by our theory reveal that multiplexing has subtle implications for threshold contagion models.
@@
We close the paper in Section 4 with observations about how multiplexing varies with individual characteristics and some implications for issues of inequality. We find that women’s networks display significantly more multiplexing than men’s networks, and that multiplexing correlates negatively with the number of connections a person has. Given our theoretical and empirical evidence showing how multiplexing can impede simple diffusions, this suggests that multiplexing could function as a channel limiting women’s exposure to information. More broadly, demographic differences in multiplexing imply that network-mediated contagions work differently in different subpopulations—a rich topic that we believe deserves further study.
@@
The literature on multiplex networks has begun to grow in the last decade (Contractor et al., 2011; Boccaletti et al., 2014; Kivela et al., 2014; Dickison et al., 2016; Bianconi, 2018). The recognition that people are involved in different types of relationships dates to some of the original works on network analysis (e.g., Simmel (1908)), and instances of the fact that different layers can serve different roles have been analyzed over time (Wasserman and Faust, 1994; Becker et al., 2020). More recent studies have shown that distinguishing between different networks and tracking their interplay can be important in understanding cooperative behavior (Atkisson et al., 2019; Cheng et al., 2021) as well as understanding play in network games and targeting policies to influence it (Walsh, 2019; Zenou and Zhou, 2024).
Our contributions to the literature on multiplex networks are threefold.
@@
First, we provide some of the first detailed statistical analyses of how multiple layers relate to each other in empirical social networks. Second, we show how different layers—as well as the level of correlation between layers—predict diffusion outcomes. This suggests that unidimensional theories of diffusion and contagion can miss important factors that determine the extent of diffusion. Third, we introduce a model and develop a new theoretical analysis of how correlation between layers impacts diffusion, which provides a basis for interpreting our empirical observations about multiplexing and diffusion.
@@
Our findings on the impact of multiplexing on diffusion can help inform a nascent and important literature on the incentives to form multiplexed networks (Billand et al., 2023; San Román, 2024). For example, a series of empirical studies on rural developing economies emphasize the role social networks play in risk-sharing arrangements (Townsend, 1994; Fafchamps and Gubert, 2007; Ambrus et al., 2014). This raises a fundamental question: If individuals primarily organize their relationships around risk-sharing and multiplex other relationships on top of the risk-sharing relationships, how might these structures affect the diffusion of new information or technologies? Both our empirical findings and theoretical results shed light on this issue.
2. Multiplexing in the Data
2.1. Two Data Sets
@@
2.1.1. The Microfinance Village Sample
The first dataset that we use, which we refer to as the “microfinance village sample,” comes from network data collected in Wave II of a study of 75 villages (Banerjee et al., 2013; Banerjee et al., 2024b). In 2012, that study obtained a complete census of the 16,476 households across these 75 villages. From 89.14% of these households, it then collected detailed socio-economic network data, described below. (This means that the study obtained information on 98.8% of the links. [^2] Given our focus on undirected graphs, we elicit a link as long as at least one of the two households on either end is sampled. With 89.14% of the households being sampled, for two arbitrary nodes $i$ and $j$, we compute P($i$ or $j$ in sample) = $0.98791-(1-0.89)^{2}=$ 0.9879.)
The researchers collected information on various types of interactions for each respondent, spanning social, financial, informational, kinship, and religious networks. The surveys asked respondents about the following types of relationships, each listed with an abbreviated label that we use from now on to refer to it:
@@ 2.1.1. The Microfinance Village Sample
2.1.2. The Diffusion RCT Sample
The second dataset that we use, which we call the “RCT village sample,” contains multiple network layers from a set of 68 different villages collected by Banerjee et al. (2019).
@@ 2.1.2. The Diffusion RCT Sample
In addition, we have the data from a randomized controlled trial (RCT) studying diffusion in these villages, which is the subject of Banerjee et al. (2019). This RCT provides cleanly identified estimates of diffusion, allowing us to examine how diffusion varies with different aspects of multiplexed networks. Specifically, in each village either 3 or 5 individuals (determined uniformly at random) were seeded with information about a promotion. Villagers could obtain a non-rivalrous chance to win either cash prizes or a mobile phone by calling in to register for the promotion. [^3] In particular, they had to dial the provided promotional number and leave a “missed call.” This was a call that we registered but did not answer and was free for the participant to make, which was a standard technique for registration at the time. Registered callers were visited a few weeks later and received a reward. [^4] The individual rolled a pair of dice and received INR 25 $\times$ the number rolled. This yielded cash prizes of amounts ranging from INR 50 (for a 2) to INR 275 (for an 11). A roll of 12 was rewarded with a cell phone worth INR 3000. The expected value of the prize was INR 255, which was more than half of a day’s wage in the area. Thus, the experiment induced the diffusion of a non-rivalrous, valuable piece of information. The outcome variable of interest is the number of households that registered. There was exogenously randomized variation in the position of the random seeds in the network, and more central seeds caused larger diffusions. We use our data on multiplexing to examine how diffusion depends on the network statistics of the seeds in various network layers, used individually and in combination, and on network multiplexing levels.
2.2. The Layers and Multiplexing Patterns
We begin with some descriptive statistics on the network layers. A link is present in a given layer if either household named the other household in one of the questions in that category (e.g., we code a kerorice link if either household reports borrowing kerosene or rice from or lending it to the other household). In terms of notation, we define a multi-layered, undirected network for each village $v$, [^5] One can also define directed networks from our data, which we comment on at several points below. Directed links open some important but tangential questions for multiplexing, which we leave for further research. for layer $\ell=1,\ldots,L$, with $g_{ij,v}^{\ell}=1$ if either household $i$ or $j$ reported having a relationship of type $\ell$. We add another layer where $i$ and $j$ are linked if they belong to the same jati. For the Microfinance Village Sample, where GPS data are available, we construct a weighted graph where the $ij$ entry is the geographic distance between the two households.
@@ 2.2. The Layers and Multiplexing Patterns
We also build a weighted and directed network whose edge weights are the sums of indicators for links in all directed layers (thus excluding jati and geography). We call this the total network. Finally, below we describe another weighted and directed aggregate network that we build from the principal component analysis.
2.2.1. Descriptive Statistics
Table 1. Descriptive Statistics
Network degree degree S.D. density triangles clustering
Microfinance villages
     social 15.296 7.841 0.079 2635.040 0.252
     kerorice 7.029 3.834 0.037 594.160 0.259
     advice 6.158 3.835 0.032 299.120 0.168
     decision 6.553 4.309 0.034 356.040 0.169
     money 8.512 5.036 0.044 681.960 0.193
     temple 1.709 1.899 0.009 52.040 0.175
     medic 6.530 3.911 0.034 369.400 0.188
     union link 75.428 32.542 0.368 314121.027 0.862
     intersect link 0.576 0.883 0.003 7.000 0.203
     jati 68.291 34.293 0.332 310150.907 1.000
RCT villages
     social 5.711 3.626 0.031 251.271 0.185
     kerorice 4.910 3.235 0.027 176.557 0.174
     advice 4.197 3.091 0.023 124.100 0.161
     decision 4.206 3.675 0.023 125.571 0.145
     union link 55.756 27.861 0.296 150771.400 0.913
     intersect link 1.812 1.829 0.010 38.871 0.229
     jati 52.633 28.599 0.279 150117.500 1.000
Our first look at the data focuses on basic descriptive statistics, presented in Table 1. The different layers exhibit significantly different patterns of connection. For example, in both datasets, the social layer is denser than the other layers and has among the highest levels of clustering. We observe a higher variance of node degrees in the decision layer than in other comparable layers (e.g., advice). [^7] In the RCT villages the social layer is significantly denser than kerorice, advice, or decision layers (p-values $0.009$, $0.000$, and $0.000$ respectively). The advice layer has significantly less variance (p-value $0.0006$) and more clustering (p-value $0.03$) than the decision layer. Similar patterns hold in the Microfinance villages.
We also observe that the microfinance villages and RCT villages differ from each other in some descriptive statistics. Microfinance villages across all network layers are denser on average and exhibit higher levels of clustering, as can be seen in Table 1. The two samples also slightly differ in terms of village size. RCT villages have 197 households on average, while microfinance villages are larger with 216 households on average.
2.2.2. Correlations among Layers
Next, we examine the correlation among layers, pictured in Figure 1.

### v2

@@ Abstract.
Social and economic networks are often multiplexed, meaning that people are connected by different types of relationships—such as borrowing goods and giving advice. We make two contributions to the study of multiplexing and the understanding of simple versus complex contagion. On the theoretical side, we introduce a model and theoretical results about diffusion in multiplex networks. We show that multiplexing impedes the spread of simple contagions, such as diseases or basic information that only require one interaction to transmit an infection. We show, however that multiplexing enhances the spread of a complex contagion when infection rates are low, but then impedes complex contagion if infection rates become high. On the empirical side, we document empirical multiplexing patterns in Indian village data. We show that relationships such as socializing, advising, helping, and lending are correlated but distinct, while commonly used proxies for networks based on ethnicity and geography are nearly uncorrelated with actual relationships. We also show that these layers and their overlap affect information diffusion in a field experiment. The advice network is the best predictor of diffusion, but combining layers improves predictions further. Villages with greater overlap between layers—more multiplexing—experience less overall diffusion. Finally, we identify differences in multiplexing by gender and connectedness. These have implications for inequality in diffusion-mediated outcomes such as access to information and adherence to norms.
@@ 1. Introduction
People maintain many different types of relationships. For example, college students’ partners in social activities overlap with, but also differ from, the people to whom they turn during times of stress or for academic collaboration (Morelli et al., 2017; Jackson et al., 2024). Villagers in south India seek advice from one set of people, borrow money from a different set, and borrow kerosene and rice from yet a third set, and these sets partly overlap. Moreover, these networks have distinct properties (e.g., link frequencies, clustering coefficients, distributions of centrality, etc.) (Banerjee et al., 2013). Importantly, even though these different network “layers” all serve different purposes, people talk with each other in each dimension and can spread disease across all layers. How does the fact that people are embedded in a number of different, overlapping networks with distinct properties affect the diffusion of information or other things that can spread via any contact?
@@
The coexistence of distinct types of relationships among the same population is known as multiplexing (see, e.g., Kivela et al. (2014)), and the interdependence of different types of relationships has been discussed since Simmel (1908). Although there are case studies of multiplexed relationships, many basic questions remain open concerning the patterns of multiplexing in social and economic relationships, as well as how multiplexing impacts diffusion.
As we show here, multiplexing turns out to be a key to understanding differences between ‘simple’ and ‘complex’ contagion. Simple contagion refers to situations where just one interaction with an infected individual is enough to transmit a disease, belief or behavior, while complex contagion refers to situations where multiple interactions/reinforcement is needed to transmit a belief or behavior (Granovetter, 1978; Centola, 2018; Beaman et al., 2021). Previous research has shown that local clustering patterns in networks are important in whether a complex contagion can be initiated (Centola, 2010). As we prove here, multiplexing determines the overall spread of a complex contagion, and in ways that can differ from how it affects the spread of a simple contagion.
In this paper, we make two main contributions, one theoretical and one empirical, about patterns of multiplexing and how it impacts diffusion.
@@
Our theoretical contribution is the development of a model of, and theoretical results on, how multiplexing determines diffusion. We first model how multiplexing affects a simple diffusion or contagion process in which a person may be infected/informed by any single infected other (called simple contagion). We introduce a definition capturing what it means for an individual to be more multiplexed in one multilayered network than another. We prove that a more multiplexed individual is less likely to become infected for any given probability of neighbor infection. Building on this result, we demonstrate that in a standard SIS (Susceptible-Infected-Susceptible) contagion model, the steady-state infection rate decreases as individuals become more multiplexed as long as transmission is not too negatively correlated across layers. These results can be summarized by saying that multiplexing impedes “simple” diffusions. We then develop a theory of how multiplexing impacts “complex” diffusion processes—in which people only become infected or adopt a new behavior/practice if they experience sufficiently many interactions with infected alters. We show that multiplexing enhances diffusion when the infection rate is low, but then impedes it when the infection rate is high. The nonmonotonicities identified by our theory reveal that multiplexing has subtle implications for complex contagion.
The intuition behind the contrast between simple and complex contagion is as follows. In simple diffusion a single infected interaction suffices to infect a node. Diversifying contacts across multiple individuals increases the probability that at least one contact is infected, compared to concentrating multiple interactions with the same person. Consequently, multiplexing impedes simple contagion, provided that transmission is not too negatively correlated across layers. Complex diffusion involves more subtle tradeoffs. Consider an individual with two links in different layers. Whether infection is more likely when these links connect to the same friend or to different friends depends on the infection’s prevalence. When infection is already widespread, the individual likely needs only one additional interaction to reach the contagion threshold. In this regime, diversifying links increases the chance of contacting an infected individual, paralleling our discussion simple diffusion, and thus less multiplexing enhances infection. Conversely, when infection rates are low, the individual more likely requires multiple additional contacts to become infected. The probability of transmission occurring on both links is higher when they connect to the same person, since only that individual need be infected, rather than requiring both of two separate contacts to be infected. Therefore, when overall infection rates are low (less contagious processes), greater multiplexing accelerates complex diffusion.
Before presenting the theoretical results, we present our other main contribution which is to document patterns of multiplexing, as well as how it impacts a diffusion empirically. We show that multiplexing varies significantly across settings, that multiple layers are differently useful in predicting a diffusion, and that it impedes diffusion in an empirical application that matches up with a simple diffusion. In particular, our empirical analysis comes in three parts.
First, using two large datasets of multiplexed networks we examine the correlation across network layers. We document both significant correlation between different network layers (types of relationships) and meaningful differences in their patterns. We show that layers of informational relationships, financial relationships, and social relationships, among others, exhibit strong correlations in a sample of 143 villages in Karnataka, India, comprising nearly 30,000 households (Banerjee et al., 2013, 2019; Banerjee et al., 2024b). At the same time, different layers display distinct patterns and differ in density and other network statistics. We also show that proxies for social relationships that are commonly used in the peer effects literature, such as geographic proximity or co-ethnicity (in our data, being members of the same jati, or subcaste), are nearly orthogonal to the other layers. This suggests that relational variables constructed based on geographic or ethnic covariates can fail to serve as accurate surrogates for actual social and economic relationships.
Second, we show that distinctions among layers are substantively important for the study of economic outcomes, specifically the diffusion of information and behaviors. Even though people have contacts and can influence each other in any network layer, we find that the different layers are differently predictive of diffusion and combine to form a nuanced overall picture. Using data from a randomized controlled trial of information diffusion, we show that some layers are more predictive of diffusion than others—with an “advice” layer standing out—and moreover that using a combination of layers yields predictions significantly better than those based on any single layer. A combination of layers also affords better predictions than using the union or intersection of layers. These findings indicate that the layers are not simply noisy observations of a latent one-dimensional network, but instead contain information richer than any one-dimensional summary. Without properly accounting for the multiplexed nature of relationships, researchers can arrive at misleading conclusions about peer effects and influence. [^1] An additional finding shows that links are most usefully viewed as multi-dimensional. We show that, although the jati layer is the least predictive of diffusion and not a good proxy for actual relationships, combining it with other layers significantly improves diffusion predictions. Thus, although jati is not a good substitute for elicited network data, it is a valuable complement. As we discuss below, using the jati variable drastically over-predicts links within jati, and under-predicts them across jati. One conjecture as to why the jati layer helps in predicting diffusion is that patterns of information passing on the network are related to jati.
Third, as another important empirical observation that helps motivate our theoretical investigation: villages that are “more multiplexed” (have more strongly correlated layers) experience less diffusion.
@@
We close the paper with observations about how multiplexing varies with individual characteristics and some implications for issues of inequality. We find that women’s networks display significantly more multiplexing than men’s networks, and that multiplexing correlates negatively with the number of connections a person has. Given our theoretical and empirical evidence showing how multiplexing can impede simple diffusions, this suggests that multiplexing could function as a channel limiting women’s exposure to information. More broadly, demographic differences in multiplexing imply that network-mediated contagions work differently in different subpopulations—a rich topic that we believe deserves further study.
Our findings are orthogonal and complementary to the research that has examined how clustering and homophily patterns affect simple and complex contagion (Granovetter, 1978; Morris, 2000; Jackson and Yariv, 2005; Centola, 2010; Golub and Jackson, 2012; Aral and Walker, 2012; Jackson and Storms, 2025; Mosleh et al., 2021). Those focus on local densities of links and community patterns that are necessary for a behavior or infection to spread from one part of a network to another, while the multiplexing examines differences in layers and correlations between those layers and how those affect overall levels of spreading. These are distinct both in the questions asked, and the reasoning and intuitions that the answers provide.
@@
The literature on multiplex networks has begun to grow in the last decade (Contractor et al., 2011; Boccaletti et al., 2014; Kivela et al., 2014; Dickison et al., 2016; Bianconi, 2018). The recognition that people are involved in different types of relationships dates to some of the original works on network analysis (e.g., Simmel (1908)), and instances of the fact that different layers can serve different roles have been analyzed over time (Wasserman and Faust, 1994; Becker et al., 2020). More recent studies have shown that distinguishing between different networks and tracking their interplay can be important in understanding cooperative behavior (Atkisson et al., 2019; Cheng et al., 2021) as well as understanding play in network games and targeting policies to influence it (Walsh, 2019; Zenou and Zhou, 2024). New (independent) work by Shi et al. (2025) also finds that network layers are differently predictive of diffusion of a behavior in a rural setting, and proposes a method of decomposing the impact of layers.
@@
Our contributions to the literature on multiplex networks are threefold. First, we provide some of the first detailed statistical analyses of how multiple layers relate to each other in empirical social networks. Second, we show how different layers—as well as the level of correlation between layers—predict diffusion outcomes. This suggests that unidimensional theories of diffusion and contagion can miss important factors that determine the extent of diffusion. Third, we introduce a model and develop a new theoretical analysis of how correlation between layers impacts diffusion, which provides a basis for interpreting our empirical observations about multiplexing and diffusion.
@@
Our findings on the impact of multiplexing on diffusion can help inform a nascent and important literature on the incentives to form multiplexed networks (Billand et al., 2023; San Román, 2024). For example, a series of empirical studies on rural developing economies emphasize the role social networks play in risk-sharing arrangements (Townsend, 1994; Fafchamps and Gubert, 2007; Ambrus et al., 2014; Cai et al., 2015). This raises a fundamental question: If individuals primarily organize their relationships around risk-sharing and multiplex other relationships on top of the risk-sharing relationships, how might these structures affect the diffusion of new information or technologies? Both our empirical findings and theoretical results shed light on this issue.
2. The Structure of Multiplexed Networks
2.1. Two Datasets of Multiplexed Networks
@@
The Microfinance Network Data
The first dataset, which we refer to as the microfinance village sample, includes 75 villages surveyed in 2012 (Banerjee et al., 2013; Banerjee et al., 2024b). That study obtained a complete census of the 16,476 households across these 75 villages. From 89.14% of these households [^2] Given our focus on undirected graphs, we elicit a link as long as at least one of the two households on either end is sampled. With 89.14% of the households being sampled, for two arbitrary nodes $i$ and $j$, we compute P($i$ or $j$ in sample) = $1-(1-0.89)^{2}=0.9879$. the researchers also collected socio-economic network data on seven types of ties:
@@
The Diffusion RCT Network Data
The second dataset, which we refer to as the RCT village sample, includes data on 68 villages collected from a separate diffusion experiment by Banerjee et al. (2019).
@@
Network Construction and Notation
A link is present in a given layer if either household named the other household in one of the questions in that category (e.g., we code a kerorice link if either household reports borrowing kerosene or rice from or lending it to the other household). In terms of notation, we define a multi-layered, undirected network for each village $v$,for layer $\ell=1,\ldots,L$, with $g_{ij,v}^{\ell}=1$ if either household $i$ or $j$ reported having a relationship of type $\ell$. We add another layer where $i$ and $j$ are linked if they belong to the same jati. For the Microfinance Village Sample, where GPS data are available, we construct a weighted graph where the $ij$ entry is the geographic distance between the two households.
@@
We also build a weighted and directed network whose edge weights are the sums of indicators for links in all directed layers (using the raw directed nomination matrices, thus excluding jati and geography). We call this the total network. Finally, below we describe another weighted and directed aggregate network that we build from the principal component analysis.
2.2. The Diffusion Randomized Controlled Trial
In the second dataset we also have the data from a randomized controlled trial (RCT) studying diffusion in these villages, which is the subject of Banerjee et al. (2019). This RCT provides cleanly identified estimates of diffusion, allowing us to examine how diffusion varies with different aspects of multiplexed networks. Specifically, in each village either 3 or 5 individuals (determined uniformly at random) were seeded with information about a promotion. Villagers could obtain a non-rivalrous chance to win either cash prizes or a mobile phone by calling in to register for the promotion.
Thus, the experiment induced the diffusion of a non-rivalrous, valuable piece of information. The outcome variable of interest is the number of households that registered. There was exogenously randomized variation in the position of the random seeds in the network, and more central seeds caused larger diffusions. We use our data on multiplexing to examine how diffusion depends on the network statistics of the seeds in various network layers, used individually and in combination, and on network multiplexing levels.
Additional details on the two datasets appear in Appendix Section A.
2.3. Correlations across Network Layers
Across both our samples, social, informational, and financial relationships often overlap—but the alignment across layers is far from complete. The social layer is denser than the other layers and has the highest level of clustering, while the decision layer exhibits higher variance of node degree than other comparable layers (e.g. advice). Jati based links are strikingly different: they form a much denser network than other layers, but serve as a poor proxy for other types of relationships—too dense, too clustered, and too homophilous to predict the other layers. Detailed descriptive statistics can be found in Supplementary Table S1.
Fig. 1 reveals three clear patterns as we examine pairwise correlations between layers.